Part a.solve 5x^2=35
Part b. find the x intercepts of f(x) = 5x^2-35

Answer by stanbon(75887)

Part a.
solve 5x^2=35
x^2=7
x = +sqrt(7) or x = -sqrt(7)
==================================

Part b. find the x intercepts of f(x) = 5x^2-35
Let y = 0 and solve for x
Answers are the same as in part a
=======================================
Cheers,
Stan H.
Reply to stanbon@comcast.net